# Wholesale Market Participation of DERA: Competitive DER Aggregation

### v4

@@ Abstract
We consider the aggregation of distributed energy resources (DERs) by a profit-seeking aggregator participating directly in wholesale market under distribution network access constraints. We propose a competitive DER aggregator (DERA) model that maximizes the DERA’s profit while ensuring each aggregated customer gains no less surplus and pays no higher energy cost than under the regulated retail tariff. The DERA participates in wholesale electricity market as virtual storage with optimized generation offers and consumption bids derived from our competitive aggregation model. Also derived are DERA’s bid curves for the distribution network access and DERA’s profitability when competing with the regulated retail tariff. We show that, with the same distribution network access, the proposed DERA’s wholesale market participation achieves the same welfare-maximizing outcome as when its customers participate directly in the wholesale market. Numerical studies compare the proposed DERA with existing methods in terms of customer surplus and DERA profit. We empirically evaluate how many DERAs can survive in the competition at long-run equilibrium, and assess the impacts of DER adoption levels and distribution network access on short-run market outcomes.
@@
distributed energy resources and aggregation, behind-the-meter distributed generation, demand-side management, net energy metering, competitive wholesale market.
@@ I Introduction
We address open problems in the direct participation of distributed energy resource aggregators (DERAs) in the wholesale electricity market operated by regional transmission organizations and independent system operators (RTOs/ISOs), under the general framework in FERC order $2$ [2]. We focus on the aggregation strategy of a profit-seeking DERA, whose industrial, commercial, and residential customers have competing service providers, such as their incumbent regulated utilities. The central theme of this work is to develop profitable and competitive aggregation strategies to attract and retain customers. By competitive aggregation, we mean that the benefits of the DERA customers must be no less than those offered by electricity provider benchmarks. An example of such a benchmark is the incumbent utility or a community choice aggregator (CCA) adopting net energy metering (NEM) policies, offering strong incentives to prosumers with behind-the-meter (BTM) DERs [3, 4, 5]. A major barrier to DERA’s entrance to direct wholesale market participation is having an aggregation strategy and a participation model to make DER aggregation a profitable venture[6].
@@ I-A Related Work
FERC Order $2$ removes regulatory barriers for DERAs to participate in wholesale capacity, energy, and ancillary service markets. Prior studies have demonstrated DERs’ capability to provide ancillary services both empirically [9, 10] and analytically [11]. In this paper, we focus on DERA participation in energy market, where aggregators submit quantity bids [12] or price-quantity bids [13, 14], and the ISO clears the market and issues dispatch signals. Our analysis centers on market efficiency, DERA profitability, and how aggregators directly control DERs [15] to follow dispatch signals. While this study focuses on energy markets, the proposed framework is also extensible to capacity and ancillary service markets.
The growing literature on DER aggregation and wholesale market participation models broadly falls into two categories. One is through a retail market design operated by a distribution system operator (DSO) [16, 17, 18], an aggregation/sharing platform [19, 20, 21], or an energy coalition such as CCA [22, 23, 24]. For the most part, these works do not consider a profit-maximizing DERA’s active participation in the wholesale market. In particular, in [16, 19], the DSO or an aggregation platform participates in the wholesale markets with the aggregated net demand (or possibly net production), treating the wholesale market as a balancing resource.
Our approach belongs to the second category of DER aggregations, where profit-seeking DERAs aggregate both generation and flexible demand resources, participating directly in the wholesale market with bid/offer curves. To ensure secure distribution network operation, DERA obeys the allocated distribution network access limits[25] (or operating envelopes [14]), rather than considering the computationally expensive network power flow constraints. Within the framework of FERC order $2$, this type of DER aggregation has the potential to improve the overall system efficiency and reliability.
@@
Most relevant to our work is the DERA’s wholesale market participation method developed by Gao et al. [12] where the authors consider a profit-seeking DERA aggregating BTM distributed generations (DGs) and offering its aggregated generation resources to the wholesale market. In particular, their approach achieves a social surplus equal to that achievable by customers’ direct participation in the competitive wholesale market. In other words, their approach achieves the highest economic efficiency for aggregating DGs. A significant difference between [12] and this paper is that we formulate a general competitive aggregation problem that includes the regulated utility. In achieving DERA’s profit maximization, our DER aggregation and market participation models are also different from [12].
The approach proposed in [12] follows the earlier work in [13] where a Stackelberg game-theoretic model is used. Both approaches assume that the DERA elicits prosumer participation with an optimized (one-part or two-part) tariff, and the prosumer responds with its quantity to be aggregated by the DERA. The real-time wholesale market price is reflected by the variable price in [7, 8, 12]. Such a variable price conveys low but volatile wholesale prices directly to customers. To protect customers from price spikes in real-time wholesale price, methods like price caps [26] were proposed.
@@ I-B Summary of Results, Contributions, and Limitations
In this paper, we considerably extend the DERA aggregation model in [1] to one that aggregates customers across multiple locations in distribution networks and incorporates security constraints on network injection and withdrawal limits. We further investigate the competitive aggregation impact on market efficiency, price stability, and long-run equilibrium.
@@
Second, we propose a virtual storage model for DERA’s wholesale market participation compatible with the practical continuous storage facility participation considered by ISOs under FERC order 841 [28, 29]. The DERA bidding curve is derived from the closed-form solution of the proposed DERA model. While the aggregation optimization explicitly involves wholesale market LMP, the virtual storage bidding curves do not require forecasting of LMP. We show that the proposed DERA wholesale market participation results in market efficiency equal to what is achievable when DERA’s customers participate directly in the wholesale market.
Finally, we derive the benefit function of DERA over distribution network injection and withdrawal access limits. DERAs compete in the distribution network access auction proposed by [25] to acquire network access and we empirically evaluate the number of surviving DERAs in the long-run competitive equilibrium. We also present a set of numerical results, comparing the surplus distribution of the proposed competitive aggregation solution with those of various alternatives, including the regulated utility. Among significant insights gained are the higher social surplus, customer surplus, and DERA surplus achievable in the proposed competitive DERA model, when compared to other alternatives.
A few remarks are warranted regarding the scope and limitation of this paper. First, the losses in distribution systems are not considered. Second, the contingency cases where DSO rejects cleared bids and offers from DERA for reliability concerns [2] are neglected. Under the access limit allocation framework proposed in [25], reliability concerns of DER aggregation are already satisfied under normal operating conditions. Lastly, although the proposed competitive aggregation offers higher benefits to DERA customers, it does so with a non-uniform payment, which might raise equity concerns.
@@
$\overline{{\bm{C}}},\underline{{\bm{C}}}$: distribution network injection and withdrawal access limits.
@@
$N$: total number of prosumers.
@@
$U$: prosumer utility function for energy consumption.
$V$: prosumer marginal utility function.
@@ I-C Paper Organization and Notations
In Sec. II, we summarize the DER aggregation model and its main interactions. The problem of competitive DER aggregation is formulated in Sec. III where we derive the optimal aggregation solution. Sec. IV and Sec. V consider DERA’s wholesale market participation and its bidding strategies in the distribution network access auction, respectively. Numerical simulations are presented in Sec. VI. Proofs are relegated in the appendix.
A list of major designated symbols is shown in Table I. The notations used here are standard. We use boldface letters for column vectors as in ${\bm{x}}=(x_{1},\ldots,x_{n})$. In particular, $\bm{1}$ is a column vector of all ones. The indicator function is denoted by $\mathbbm{1}\{x_{n}\leq y_{n}\}$, which equals 1 if $x_{n}\leq y_{n}$, and 0 otherwise. ${\bm{x}}\preceq{\bm{y}}$ means $x_{n}\leq y_{n},\forall n$. $\mathbb{R}_{+}$ represents the set of all nonnegative real numbers. $[x]$ represents the set of integers from 1 to $x$, i.e., $[x]:=\{1,\ldots,x\}$.
@@ II DER Aggregation Model
A DERA aggregates flexible resources from its customers and coordinates with the DSO for power delivery to the wholesale market operated by ISO/RTO. Following the DERA interaction model proposed in [25], we focus on the DERA-DSO-ISO/RTO interfaces (a)–(c), as shown in Fig. 1. Since a DERA uses DSO’s physical infrastructure for power delivery between its customers and the wholesale market, it is essential to delineate the financial and physical interactions at these interfaces. Below, we describe the three interfaces (a)–(c).
@@
DERA and its customers at interface (a): We assume the DERA aggregates resources in a retail market from residential, commercial, and industrial customers, who have the option of being served by a regulated utility. Under a single-bill payment model, each customer settles both consumption charges and production compensation with the DERA. The DERA deploys an energy management system with direct controls [^2] The direct control in [15] is implemented through cloud-based platforms. of the customer’s BTM generation and flexible demand resources, such as rooftop PV, heat pumps, water heaters, and EV chargers. To remain competitive with the incumbent utility, the DERA must offer more attractive aggregation benefits; otherwise, customers will revert to the utility. Accordingly, the DERA optimizes the customer’s BTM resources and provides the customer with a cost-benefit comparison with the NEM benchmark offered by the utility. See Sec. III for details.
@@ III Competitive DER Aggregation
This section formulates the optimal competitive aggregation and analyzes the properties of the optimal solution when competing with the incumbent utility’s NEM. Our DER aggregation is built on the deregulated retail market. For example, in Texas and New York, [^3] For reference, we provide links to publicly available listings of retail energy providers in Texas and New York. customers can choose their electricity suppliers based on electricity rate and services. We consider prosumers owning all energy consumption and DG devices. After joining a DERA, prosumers grant device access to DERA for measurements and direct controls.
@@ III-A Closed-Form Solution for Competitive DER Aggregation
We consider a DERA aggregating customers over multiple points of aggregation (PoAs) in the distribution network. [^4] For simplicity, we illustrate the single time interval aggregation model here and apply it to the multi-interval aggregation empirically in Sec. VI-E. We define PoAs [^5] A diagram illustrating PoA is in Fig. 2 of [25]. as the main buses with higher voltages in the distribution network, which can be recognized with main substation information [30]. With the DERA-DSO coordination method in [25], [^6] Details about how the DERA coordinates with DSO to get distribution network injection and withdrawal accesses are explained in Sec. V. DERA receives injection and withdrawal access limits at all PoAs, represented respectively by
@@
where $\overline{{\bm{C}}},\underline{{\bm{C}}}\in\mathbb{R}_{+}^{M}$ and $M$ denotes the total number of PoAs. Thus, in the real-time operation, DERA must guarantee that its aggregated power at the $m$-th PoA satisfies
$-\underline{C}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}(g_{n}-\mathbf{1}^{\top}{\bm{d}}_{n})\leq\overline{C}_{m},\forall m\in[M],$ (1)
@@
$U_{n}({\bm{d}}_{n})$ is the $n$-th customer’s utility of consuming ${\bm{d}}_{n}$. We assume the utility function is concave, nonnegative, nondecreasing, continuously differentiable, additive (i.e., across the $K$ devices $U({\bm{d}})=\sum_{k=1}^{K}U_{k}(d_{k})$). Here, the utility function is given; in practice, utility functions can be computed by parametric [27] or nonparametric [31] methods.
${\cal K}$-competitive constraint is the criterion for a rational customer, seeking surplus maximization, to join a DERA. Otherwise, a rational prosumer has the incentive to leave DERA and switch to the benchmark service provider for a higher customer surplus. More details about the benchmark prosumer surplus ${\cal K}_{n}$ are explained in Sec. III-B.
[…]
${\bm{D}}:=(\bm{d}_{n},n\in[N]),~{}~{}\hbox{\boldmath$\omega$\unboldmath}:=(\omega_{n},n\in[N])$
@@
The optimal value $\Pi(\overline{{\bm{C}}},\underline{{\bm{C}}})$ is DERA’s profit given the distribution network access. In the objective function, DERA seeks profit maximization from both the aggregated customers’ payments and the revenue from the wholesale market. Without loss of generality, all PoAs are under the same point of interconnection, facing a common LMP $\pi\in\mathbb{R}$.
Under the above assumptions for feasibility in (4) and the utility function, Theorem 1 establishes a closed-form solution of (6) parameterized by $\pi$.
@@
where $V(x):=\frac{d}{dx}U(x)$ is the marginal utility function and
$h_{nk}(x):=\max\{\underline{d}_{nk},\min\{V_{nk}^{-1}(x),\overline{d}_{nk}\}\}.$ (9)
$m$ is the index of PoA connecting customer $n$, i.e., $n\in{\cal N}_{m}$ defined by (2). Solve for $\underline{\xi}_{m},\overline{\xi}_{m},\forall m\in[M]$ from
@@
The proof is in Appendix VIII-A, following the convexity and Karush-Kuhn-Tucker (KKT) conditions of (6). This optimal solution has two noteworthy characteristics. First, the optimal consumption in (7) is only a function of the LMP $\pi$ when DERA purchases enough network accesses at the PoA. Note also the difference between the optimal consumption schedule in (7) and those in [13] where the optimally scheduled consumption always depends on the anticipated LMP and forecast of BTM DG. Second, (6) finds a Pareto efficient allocation that maximizes the surplus of the DERA, subject to the constraint that the aggregated customer has the given level of surplus ${\cal K}_{n}$. Similar optimization and the Pareto efficient allocation are also analyzed in [32, P602] for first-degree price discrimination. The payment function $\omega_{n}$ can be realized by a two-part tariff, which is explored in [12], although it unidirectionally aggregates BTM DG. Overall, such a closed-form solution allows DERA to apply simple dispatch and pricing policies when aggregating massive numbers of households over multiple PoAs in the distribution networks.
III-B Properties of DERA Competitive with NEM X
We analyze the profitability of DERA and the energy consumption cost of aggregated prosumers when DERA is competitive with a regulated NEM tariff parameterized by the retail (consumption) rate $\pi^{+}$, the sell (production) rate $\pi^{-}$, and the connection charge $\pi^{0}$. Assume $0\leq\pi^{-}\leq\pi^{+}$ and customers’ surpluses under NEM are nonnegative [27]. Denote the $n$-th prosumer surplus under NEM to be $S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n})$, whose computation depends on the DG generation and network access limits. Explicit formulations are in Sec. VIII-B. DERA sets the benchmark prosumer surplus

### v5

# Wholesale Market Participation via Competitive DER Aggregation

@@ Abstract
We consider the aggregation of distributed energy resources (DERs), such as solar PV, energy storage, and flexible loads, by a profit-seeking aggregator participating directly in the wholesale market under distribution network access constraints. We propose a competitive DER aggregator (DERA) model that directly controls local DERs to maximize its profits, while ensuring each aggregated customer gains a surplus higher than their surplus under the regulated retail tariff. The DERA participates in the wholesale electricity market as virtual storage with optimized generation offers and consumption bids derived from the propoed competitive aggregation model. Also derived are DERA’s bid curves for the distribution network access and DERA’s profitability when competing with the regulated retail tariff. We show that, with the same distribution network access, the proposed DERA’s wholesale market participation achieves the same welfare-maximizing outcome as when its customers participate directly in the wholesale market. Extensive numerical studies compare the proposed DERA with existing methods in terms of customer surplus and DERA profit. We empirically evaluate how many DERAs can survive in the competition at long-run equilibrium, and assess the impacts of DER adoption levels and distribution network access on short-run operations.
@@
distributed energy resources and aggregation, demand-side management, direct control, distribution network, net energy metering, competitive wholesale market.
@@ I Introduction
We address open problems in the direct participation of distributed energy resource aggregators (DERAs) in the wholesale electricity market operated by regional transmission organizations and independent system operators (RTOs/ISOs), as envisioned by FERC order 2222 [2]. We focus on the aggregation strategy of a profit-seeking DERA, whose industrial, commercial, and residential customers have competing service providers, such as their incumbent regulated utilities. Our main objective is to design aggregation strategies that allow the DERA to remain profitable while offering competitive aggregation services that attract customers and grant the DERA operational control of their DERs. [^1] Direct DERA control over DERs is inspired by programs such as the Tesla Virtual Power Plant, which establishes a control protocol once customers opt into the aggregation service [3]. By competitive aggregation, we mean that the benefits of the DERA customers must be no less than those offered by electricity provider benchmarks. An example of such a benchmark is the incumbent utility or a community choice aggregator (CCA) adopting net energy metering (NEM) tariffs, offering strong incentives to prosumers with behind-the-meter (BTM) DERs [4, 5, 6]. A major barrier to DERA’s entrance to direct wholesale market participation is having an aggregation strategy and a participation model to make DER aggregation a profitable venture, while competitively attracting customers [7].
@@ I-A Related Work
FERC Order 2222 removes regulatory barriers for DERAs to participate in wholesale capacity, energy, and ancillary service markets. In this paper, we focus on DERA participation in energy market, where aggregators submit quantity [10] or price-quantity bids [11], and the ISO clears the market and issues dispatch signals. Our analysis centers on market efficiency, DERA profitability, and how aggregators directly control DERs [3, 12] to follow dispatch signals. While this study focuses on energy markets, the proposed framework is also extensible to capacity and ancillary service markets.
The growing literature on DER aggregation and wholesale market participation models broadly falls into two categories. One is through a distribution network optimization operated by a distribution system operator (DSO) [13, 14], an aggregation/sharing platform [15, 16], or an energy coalition [17, 18]. For the most part, these works do not consider a profit-maximizing DERA’s active participation in the wholesale market at the transmission grid. In particular, in [13, 15, 14], the DSO or an aggregation platform participates in the wholesale markets with the aggregated power, treating the transmission network and wholesale market as a balancing resource.
Our approach belongs to the second category of DER aggregations, where profit-seeking DERAs aggregate both generation and flexible demand resources, participating directly in the wholesale market with bid/offer curves. To ensure secure distribution network operation, DERA obeys the allocated distribution network access limits [19] (a.k.a. operating envelopes [18] or feasibility sets [20, 21]), rather than considering the computationally expensive network power flow constraints. With the direct wholesale market participation supported by FERC Order 2222, this type of DER aggregation has the potential to improve the overall system efficiency and reliability.
@@
Most relevant to our work is the DERA’s wholesale market participation method developed by Gao et al. [10] where the authors consider a profit-seeking DERA aggregating BTM distributed generations (DGs) and offering its aggregated generation resources to the wholesale market. In particular, their approach achieves a social surplus equal to that achievable by customers’ direct participation in the competitive wholesale market. In other words, their approach achieves the highest economic efficiency for aggregating DGs. A significant difference between [10] and this paper is that we formulate a general competitive aggregation that includes the regulated utility. In achieving DERA’s profit maximization, our aggregation and market participation are also different from [10].
The approach proposed in [10] follows the earlier work in [11] where a Stackelberg game-theoretic model is used. Both approaches assume that the DERA elicits prosumer participation with an optimized (one-part or two-part) tariff, and the prosumer responds with its quantity to be aggregated by the DERA. The real-time wholesale market price is reflected by the variable price in [8, 9, 10]. Such a variable price conveys low but volatile wholesale prices directly to customers. To protect customers from price spikes in real-time wholesale prices, methods like price caps [22] were proposed.
@@ I-B Summary of Results, Contributions, and Limitations
In this paper, we substantially extend the DERA aggregation model in [1] to one that controls aggregated customers across multiple locations in distribution networks and incorporates security constraints on network injection and withdrawal limits. We further investigate the competitive aggregation impact on market efficiency, price stability, and long-run equilibrium.
@@
Second, we propose a virtual storage model for DERA’s wholesale market participation compatible with the practical continuous storage facility participation considered by ISOs [23, 24]. The DERA bidding curve is derived from the closed-form solution of the proposed DERA model. While the aggregation optimization explicitly involves wholesale market LMP, the virtual storage bidding curves do not require forecasting of LMP. We show that the proposed DERA wholesale market participation results in market efficiency equal to what is achievable when DERA’s customers participate directly in the wholesale market.
Finally, we derive the benefit function of DERA over distribution network injection and withdrawal access limits. DERAs compete in the distribution network access auction proposed by [19] to acquire network access, and we empirically evaluate the number of surviving DERAs in the long-run competitive equilibrium. We also present a set of numerical results, comparing the surplus distribution of the proposed competitive aggregation solution with those of various alternatives, including the regulated utility. Among significant insights gained are the higher social surplus, customer surplus, and DERA surplus achievable in the proposed competitive DERA model, when compared to other alternatives.
A few remarks are warranted regarding the scope and limitations of this paper. First, the losses in distribution systems are not considered. Second, the contingency cases where DSO rejects cleared bids and offers from DERA for reliability concerns [2] are neglected. Under the access limit allocation framework proposed in [19], reliability concerns of DER aggregation are already satisfied under normal operating conditions. Lastly, although the proposed competitive aggregation offers higher benefits to DERA customers, it does so with a non-uniform payment, which might raise equity concerns.
@@
$\overline{{\bm{C}}},\underline{{\bm{C}}}$: distribution network injection and withdrawal limits.
@@
NN: total number of prosumers.
@@
$U(\cdot)$: prosumer utility of energy consumption function.
$V(\cdot)$: prosumer marginal utility function.
@@ I-C Paper Organization and Notations
In Sec. II, we summarize the DER aggregation model and its main interactions. The problem of competitive DER aggregation is formulated in Sec. III where we derive the optimal aggregation solution. Sec. IV and Sec. V consider DERA’s wholesale market participation and its bidding strategies in the distribution network access auction, respectively. Numerical simulations are presented in Sec. VI. Proofs for some of the major theoretical results are presented in the appendix.
A list of major designated symbols is shown in Table I. The notations used here are standard. We use boldface letters for column vectors as in ${\bm{x}}=(x_{1},\ldots,x_{n})$. In particular, $\bm{1}$ is a column vector of all ones. The indicator function is denoted by $\mathbbm{1}\{x_{n}\leq y_{n}\}$, which equals 1 if $x_{n}\leq y_{n}$, and 0 otherwise. ${\bm{x}}\preceq{\bm{y}}$ means $x_{n}\leq y_{n},\forall n$. $\mathbb{R}_{+}$ represents the set of all nonnegative real numbers. $[x]$ represents the set of integers from 1 to $x$, i.e., $x]:=\{1,\ldots,x\}$.
@@ II DER Aggregation Model
A DERA aggregates flexible resources from its customers and coordinates with the DSO for power delivery to the wholesale market operated by ISO/RTO. Following the DERA interaction model proposed in [19], we focus on the DERA-DSO-ISO/RTO interfaces (a)–(c), as shown in Fig. 1. Since a DERA uses DSO’s physical networks for power delivery between its customers and the wholesale market, it is essential to delineate the financial and physical interactions at these interfaces. Below, we describe the three interfaces (a)–(c).
@@
DERA and its customers at interface (a): We assume the DERA aggregates BTM resources from retail market customers in the distribution network, where each customer alternatively has the option of being served by a regulated utility. Under a single-bill payment model, each customer settles both consumption charges and production compensation with the DERA. The DERA deploys an energy management system that controls customer’s BTM generation and flexible demand resources, such as rooftop PV, heat pumps, water heaters, and EV chargers. To remain competitive with the incumbent utility, the DERA must offer more attractive aggregation benefits; otherwise, customers will revert to the utility. These competitive aggregation benefits motivate customers to opt into the aggregation program, which includes an agreement granting the DERA direct control of their DERs. Customers retain the ability to operate their own DERs when desired, but the control agreement enables the DERA to coordinate DERs when participating in the wholesale market [3]. Accordingly, the DERA optimizes and dispatches customers’ BTM resources and provides each customer with a cost–benefit comparison relative to the NEM benchmark offered by the utility. See Sec. III for details.
@@ III Competitive DER Aggregation
This section formulates the optimal competitive aggregation and analyzes the properties of the optimal solution when competing with the incumbent utility’s NEM. Our DER aggregation is built on the deregulated retail market. For example, in Texas [25] and New York [26], customers can choose their electricity suppliers based on electricity rate and services. We consider prosumers owning all energy consumption and DG devices. After joining a DERA, prosumers grant device access to the DERA for measurements and direct controls.
@@ III-A Closed-Form Solution for Competitive DER Aggregation
We consider a DERA aggregating customers over multiple points of aggregation (PoAs) in the distribution network. [^3] For simplicity, we illustrate the single time interval aggregation model here and apply it to the multi-interval aggregation empirically in Sec. VI-E. We define PoAs as the main buses with higher voltages in the distribution network, which can be recognized with main substation information [27]. A diagram illustrating PoA is in Fig. 2 of [19]. With the DERA-DSO coordination method in [19], DERA receives injection and withdrawal access limits at all PoAs, represented respectively by
@@
where $\overline{{\bm{C}}},\underline{{\bm{C}}}\in\mathbb{R}_{+}^{M}$ and $M$ denotes the total number of PoAs. Details about how the DERA coordinates with DSO to get distribution network injection and withdrawal accesses are explained later in Sec. V. Thus, in the real-time operation, DERA must guarantee that its aggregated power at the $m$-th PoA satisfies
$\underline{C}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}(g_{n}-\mathbf{1}^{\top}{\bm{d}}_{n})\leq\overline{C}_{m},\forall m\in[M],$ (1)
@@
$U_{n}({\bm{d}}_{n})$ is the $n$-th customer’s utility of consuming ${\bm{d}}_{n}$. We assume the utility function is concave, nonnegative, nondecreasing, continuously differentiable, and additive (i.e., across the $K$ devices $U({\bm{d}})=\sum_{k=1}^{K}U_{k}(d_{k})$). Here, the utility function is given; in practice, utility functions can be computed by parametric [5] or nonparametric [28] methods.
$\cal K$}-competitive constraint is the criterion for a rational customer, seeking surplus maximization, to join a DERA. Otherwise, a rational prosumer has the incentive to leave DERA and switch to the benchmark service provider for a higher customer surplus. More details about the benchmark prosumer surplus ${\cal K}_{n}$ are explained in Sec. III-C.
To summarize, in real-time, the DERA solves for the consumption bundle of all aggregated customers ${\bm{D}}\in\mathbb{R}_{+}^{N\times K}$ and their single-bill payments $\hbox{$\omega$}\in\mathbb{R}^{N}$, defined by
${\bm{D}}:=(\bm{d}_{n},n\in[N]),~~\hbox{$\omega$}:=(\omega_{n},n\in[N])$
@@
The optimal value $\Pi(\overline{{\bm{C}}},\underline{{\bm{C}}})$ is DERA’s profit given the distribution network access. In the objective function, DERA seeks profit maximization from both the aggregated customers’ payments and the revenue from the wholesale market. Without loss of generality, all PoAs are under the same point of interconnection, facing a common LMP $\pi\in\mathbb{R}$. Optimal solutions and values in (6) support the control of aggregated DERs, offer/bid curves to the wholesale market (Sec. IV), and also the distribution network access request to DSO (Sec.V).
Under the above assumptions for feasibility in (4) and the utility function, Theorem 1 below establishes a closed-form solution of (6) parameterized by $\pi$. We denote $V(x):=\frac{d}{dx}U(x)$ as the marginal utility function, define
$h_{nk}(x):=\max\{\underline{d}_{nk},\min\{V_{nk}^{-1}(x),\overline{d}_{nk}\}\},$ (7)
and solve for $\underline{\xi}_{m},\overline{\xi}_{m},\forall m\in[M]$ from
@@
where $m$ is the index of PoA connecting customer $n$, i.e., $n\in{\cal N}_{m}$ defined by (2).
@@
The proof is in the appendix, following the convexity and Karush-Kuhn-Tucker (KKT) conditions of (6). This optimal solution has two noteworthy characteristics. First, the optimal consumption in (10) is only a function of the LMP $\pi$ when DERA purchases enough network accesses at the PoA. Note also the difference between the optimal consumption schedule in (10) and those in [11], where the optimally scheduled consumption always depends on the anticipated LMP and forecast of BTM DG. Second, (6) finds a Pareto efficient allocation that maximizes the surplus of the DERA, subject to the constraint that the aggregated customer has the given level of surplus ${\cal K}_{n}$. Similar optimization and the Pareto efficient allocation are also analyzed in [29, P602] for first-degree price discrimination. The payment function $\omega_{n}$ can be realized by a two-part tariff, which is explored in [10], although it unidirectionally aggregates BTM DG. Overall, such a closed-form solution allows DERA to apply simple dispatch and pricing policies when aggregating massive numbers of households over multiple PoAs in the distribution networks.
III-B NEM benchmarks
Considering the benchmark performance of a regulated utility offering the NEM tariff, we extend the results in [5] and present closed-form characterizations of consumer/prosumer surpluses. Detailed step-by-step derivations are in Sec. VIII-B of the appendix. For simplicity, we consider one representative prosumer by setting ${\cal N}_{m}=1,K=1$ and dropping the prosumer index $n$ and PoA index $m$. The prosumer’s net consumption is z:=dg,z:=d-g, where $g\in[0,\infty)$ is the BTM DG. The prosumer is a producer if $z<0$ and a consumer if $z\geq 0$.
In evaluating the benchmark prosumer surplus under a regulated utility, we assume that the prosumer maximizes its surplus under the utility’s NEM tariff, where $\pi^{+}$ is the retail (consumption) rate, $\pi^{-}$ the sell (production) rate, and $\pi^{0}$ the connection charge. In general $\pi^{-}\leq\pi^{+}$ under NEM tariff, and the prosumer’s energy bill $P(z)$ for the net consumption $z$ is given by the convex function $P(z):=\max\{\pi^{+}z,\pi^{-}z\}+\pi^{\mbox{\tiny 0}}.$ The prosumer surplus under NEM is S(d):=U(d)P(z).S(d):=U(d)-P(z).
For an active prosumer whose consumption is a function of the available DG output $g$, the optimal consumption can be obtained by $d_{\mbox{\tiny NEM-a}}:=\arg\max_{d\in{\cal D}}(U(d)-P(d-g)).$ For the fairness of comparison, we assume the aggregated customer is subject to the same distribution network injection and withdrawal access limits, i.e., $\underline{C}\leq g-d\leq\overline{C}$, which is the same as that applied to the proposed DERA optimization (6). So, for the above optimization, the domain is ${\cal D}:=[\max\{\underline{d},g-\overline{C}\},\min\{\bar{d},g+\underline{C}\}]$.
The surplus $S_{\mbox{\tiny NEM-a}}$ and the consumption $d_{\mbox{\tiny NEM-a}}$ of an active prosumer are given by the following equations.
$S_{\mbox{\tiny NEM-a}}(g,\underline{C},\overline{C})=U(d_{\mbox{\tiny NEM-a}})-P(d_{\mbox{\tiny NEM-a}}-g)$ (12)
[…]
$d_{\mbox{\tiny NEM-a}}=\max\{d^{+},\min\{g,d^{-}\}\},$
where $d^{+}:=f(\pi^{+})$, $d^{-}:=f(\pi^{-})$, $d^{0}:=f(\mu^{*}(g))$ with
$f(x):=\max\{\underline{d},g-\overline{C},\min\{V^{-1}(x),\bar{d},g+\underline{C}\}\},$ (13)
and, by solving $f(\mu)=g$, we have $\mu^{*}(g)\in[\pi^{-},\pi^{+}]$.
A prosumer is called passive if it decides energy consumption without the awareness of its DG output and the influence brought by NEM X switching among $\pi^{-}$ and $\pi^{+}$. The optimal consumption bundle of such a passive prosumer under the NEM X tariff is given by $d_{\mbox{\tiny NEM-p}}:=\arg\max_{d\in{\cal D}}(U(d)-\pi^{+}d).$ The total consumption $d_{\mbox{\tiny NEM-p}}$ and the surplus $S_{\mbox{\tiny NEM-p}}$ of a passive prosumer are given by
$S_{\mbox{\tiny NEM-p}}(g,\underline{C},\overline{C})=U(d_{\mbox{\tiny NEM-p}})-P(d_{\mbox{\tiny NEM-p}}-g)$ (14)
$=\begin{cases}U(d^{+})-\pi^{-}(d^{+}-g)-\pi^{0},&g\geq d^{+}\\ U(d^{+})-\pi^{+}(d^{+}-g)-\pi^{0},&g<d^{+}\end{cases}$
$d_{\mbox{\tiny NEM-p}}=d^{+}.$ (15)